# Large deviations for the maximum of a reducible two-type branching Brownian motion

### v1

@@ Abstract.
We consider a two-type reducible branching Brownian motion, defined as a particle system on the real line in which particles of two types move according to independent Brownian motions and create offspring at a constant rate. Particles of type $1$ can give birth to particles of types $1$ and $2$, but particles of type $2$ only give birth to descendants of type $2$. Under some specific conditions, Belloum and Mallein in [4] show that the maximum position $M_{t}$ of all particles alive at time $t$, suitably centred by a deterministic function $m_{t}$, converge weakly. In this work, we are interested in the decay rate of the following upper large deviation probability, as $t\rightarrow\infty$,
@@ 1. Introduction and main results
In this article, we study the large deviation probabilities for the maximum of the two-type reducible BBM. That is the convergence rate of the following probability
@@
If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$, then the following limit exists for any $\theta>1$,
@@
where for $\beta<1$,
@@
for $\beta>1$ and $\sigma^{2}\geq 1$, we have
$A=\beta\left(1-\theta^{2}\right),$
for $\beta>1$ and $\sigma^{2}<1$, we have
$A=\begin{cases}\beta\left(1-\theta^{2}\right),&\theta^{2}\leq\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta},\\ 1-\theta^{2}\beta\sigma^{2},&\theta^{2}\geq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta},\\ \frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}\leq\theta^{2}\leq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.\end{cases}$
@@
If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$, then the following limits exists for any $\theta>1$,
@@
where for $\beta>1$, we have
@@
for $\beta\leq 1,\sigma^{2}\leq 1$, we have
$A=1-\theta^{2},$
for $\beta\leq 1,\sigma^{2}>1$, we have
@@
If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$, then for any $\theta>1$, we have
@@
where $C(\theta,\bar{u})$ is a constant depending only on $\theta$ and $\bar{u}$; see [11], [2] and [5]. In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of two-type inreduciable BBM. It is possible to obtain more preciser form as (3) with the help of truncated second moment method and the F-KPP equation. Our method may also apply to study the cases of $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{II}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}}\cap\mathcal{C}_{\text{III}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{III}}$; see [17] for related results.
@@ 3.1. Upper bound
i). If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\leq 1$, then $u\mapsto f_{1}^{\prime}(u)$ is an increasing function. In this case, by noticing that $\beta\sigma^{2}>1$, one has
@@
ii). If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\geq 1,\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
@@
iii). If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
@@
First, from 3, we have
@@
Solving $f^{\prime}(u_{0})=0$, we have
@@
First, from 3, we have
@@
This corresponds to the cases that $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, or $\beta\leq 1,\sigma^{2}>1$ and $\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}.$ The arguments are parallel to those in Section 3.2.1. We only give an outline. Choose
@@
This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Again, we only give an outline of the proof. Choose
@@ 5.1. Upper bound
Solving $f^{\prime}(u_{0})=0$ yields
@@
First, from 3, we have

### v2

@@ Abstract.
We consider a two-type reducible branching Brownian motion, defined as a particle system on the real line in which particles of two types move according to independent Brownian motions and create offspring at a constant rate. Particles of type $1$ can give birth to particles of types $1$ and $2$, but particles of type $2$ only give birth to descendants of type $2$. Under some specific conditions, Belloum and Mallein in [4] showed that the maximum position $M_{t}$ of all particles alive at time $t$, suitably centered by a deterministic function $m_{t}$, converge weakly. In this work, we are interested in the decay rate of the following upper large deviation probability, as $t\rightarrow\infty$,
@@ 1. Introduction and main results
In this article, we study the large deviation probabilities for the maximum of a two-type reducible BBM. That is the convergence rate of the following probability
@@
Assume $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$. Then the following limit exists for any $\theta>1$,
@@
where if $\beta\leq 1$, or $\beta>1$ and $\sigma^{2}\geq 1$, then
@@
and if $\beta>1$ and $\sigma^{2}<1$, then we have
$A=\begin{cases}\beta\left(1-\theta^{2}\right),&\theta^{2}\leq\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta},\\ 1-\theta^{2}\beta\sigma^{2},&\theta^{2}\geq\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta},\\ \frac{-2\sqrt{\beta(\beta-1)\left(1-\sigma^{2}\right)}\sigma\theta+\beta-\sigma^{2}}{1-\sigma^{2}},&\frac{(\beta-1)\sigma^{2}}{\left(1-\sigma^{2}\right)\beta}<\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)\sigma^{2}\beta}.\end{cases}$
@@
Assume $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{II}}$. Then the following limit exists for any $\theta>1$,
@@
where if $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, we have
@@
and if $\beta\leq 1,\sigma^{2}>1$, we have
@@
Assume $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{III}}$. Then for any $\theta>1$, we have
@@
where $C(\theta,\bar{u})$ is a constant depending only on $\theta$ and $\bar{u}$; see [11], [2] and [5]. In this work, we only obtain the exponential decay rates for the large deviation probabilities for the maximum of a two-type inreduciable BBM. It is possible to obtain a form as (3) with the help of the so-called truncated second moment method and the F-KPP equation. Our method may also apply to study the cases of $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{II}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{II}}\cap\mathcal{C}_{\text{III}}$ or $(\beta,\sigma^{2})\in\mathcal{C}_{\text{I}}\cap\mathcal{C}_{\text{III}}$; see [17] for related results.
@@ 3.1. Upper bound
Case I. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\leq 1$, then $u\mapsto f_{1}^{\prime}(u)$ is an increasing function. In this case, by noticing that $\beta\sigma^{2}>1$, one has
@@
Case II. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta\geq 1,\sigma^{2}\geq 1$, then $f_{1}^{\prime}(u)\geq 0$. Thus $\max_{0\leq u\leq 1}f_{1}(u)=f_{1}(1)$ and hence
@@
Case III. If $\left(\beta,\sigma^{2}\right)\in\mathcal{C}_{\text{I}}$ and $\beta>1,\sigma^{2}<1$, then obviously $u\mapsto f_{1}^{\prime}(u)$ is decreasing. Then we have three subcases.
@@
First, from (3), we have
@@
Solving $f_{2}^{\prime}(u_{0})=0$, we have
@@
First, from (3), we have
@@
This corresponds to the cases that $\beta>1$, or $\beta\leq 1,\sigma^{2}\leq 1$, or $\beta\leq 1,\sigma^{2}>1$ and $\theta^{2}<\frac{(\beta-1)}{\left(1-\sigma^{2}\right)}.$ Choose
@@
This case corresponds to the case that $\beta\leq 1,1<\sigma^{2}\leq\frac{1}{\beta}$ and $\theta^{2}\geq\frac{(1-\beta)}{\left(\sigma^{2}-1\right)}\sigma^{4}.$ Choose
@@ 5.1. Upper bound
Solving $f_{3}^{\prime}(u_{0})=0$ yields
@@
First, from (3), we have